# Machine learning Hubbard parameters with equivariant neural networks

### v1

@@ DISCUSSION
The usage of our ML model is straightforward and entails two DFT-based calculations for a given material. Initially, a ground-state calculation employing DFT+$U$+$V$ with initial guesses for $U$ and $V$ (which can be set to zero) is conducted to determine the atomic occupation matrices required as input for the model. Subsequently, a final structural optimization using the model-predicted SC Hubbard parameters yields a self-consistent structural-electronic ground state. The computational cost of these two calculations and a model evaluation is negligible compared to the DFPT-based SC protocol they replace. Furthermore, we have demonstrated that the full SC protocol can be reduced to just a few iterations, which prove adequate for achieving a RMSE within a few percent of the real value.
@@
Unlike other ML models designed for predicting Hubbard parameters, [28, 61, 29, 62] our model does not rely on experimental data or information from other state-of-the-art computational methods such as $GW$ and hybrid functionals, which possess inherent limitations and specific ranges of applicability. Instead, our model exclusively relies on linear-response theory through DFPT, which provides material-specific Hubbard parameters directly reflecting the local chemistry and OS of TM elements. Notably, our model not only predicts on-site Hubbard $U$ parameters but also inter-site Hubbard $V$ parameters, crucial for materials characterized by significant covalent interactions, an aspect that was disregarded in previous studies. [81] Furthermore, the architecture of our ML model is highly versatile, permitting easy integration of additional inputs and outputs, facilitating exploration of diverse learning tasks beyond learning Hubbard parameters. We expect that similar hybrid ML-accelerated electronic structure methodologies, maintaining accuracy and transferability, will become prevalent, potentially yielding a comparable impact on the field as observed with ML interaction potentials.
@@
Lastly, it is essential to highlight the limitations of the trained models we have presented. These were trained on data generated using a specific computational setup (see below), which must remain exactly the same when the model is applied to other systems. In other words, the $U$ and $V$ values predicted by the model are not transferable across electronic-structure codes and even across different pseudopotentials within the same code. [82] Furthermore, the training data only encompass Fe, Mn, and Ni with Hubbard $U$ corrections, while $V$ parameters are available for pairs involving these elements and O. Consequently, these models can effectively predict Hubbard parameters for other materials with Hubbard corrections on these atoms or atom pairs, provided that the model input is generated using identical pseudopotentials, Hubbard projectors, and functional (further elaborated in the methods section). Nevertheless, with a more diverse dataset, the models could be easily extended to accommodate a broader range of compositions. Indeed, ongoing efforts aim to establish a comprehensive database of Hubbard parameters for various TM-containing materials, akin to Ref. 83, thereby significantly broadening the scope of our model to encompass numerous TM elements across diverse OS and chemical environments. Moreover, our model can integrate into automated AiiDA workflows, [84, 85] enabling non-experts to harness it effortlessly and access SC Hubbard parameters with minimal intervention. Consequently, we believe that our ML model represents a significant advancement in expediting materials discovery, design, and understanding based on the DFT+$U$+$V$ approach, thereby unlocking new avenues for technological progress and breakthroughs.
@@ METHODS
All calculations are performed using the plane-wave pseudopotential method as implemented in the Quantum ESPRESSO  distribution. [86, 87, 88] We use the exchange-correlation functional constructed using $\sigma$-GGA with the PBEsol prescription, [89] and the pseudopotentials are taken from the SSSP library v1.1 (efficiency). [90, 91] For metallic ground states, we use Gaussian smearing. To construct the Hubbard projectors, we use atomic orbitals which are orthonormalized using Löwdin’s method. [92, 93] Structural optimizations are performed using DFT+$U$+$V$ [58] with the Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm [94] and convergence thresholds for the total energy of $10^{-6}$ Ry, for forces of $10^{-5}$ Ry/Bohr, and for pressure of $0.5$ Kbar. The DFPT calculations of Hubbard parameters are performed using the HP code, [95] with an accuracy of $0.01-0.1$ eV for the computed values of $U$ and $V$. The information about the kinetic-energy cutoff, $\mathbf{k}$ and $\mathbf{q}$ points sampling of the Brillouin zone for each system are detailed in Sec. S9 in the SI. More technical details can be found directly in the source files publicly available through the Materials Cloud Archive (see below).
@@
The data used to produce the results of this work will be available in the Materials Cloud Archive prior to the publication at https://archive.materialscloud.org.
@@
The code for the equivariant neural networks is available at https://github.com/camml-lab/hubbardml/. Results were generated using the code at git commit d67007e04676741818f5bcebb4373746351b4b3d.

### v2

@@ DISCUSSION
The usage of our ML model is straightforward and entails two DFT-based calculations for a given material. Initially, a ground-state calculation employing DFT+$U$+$V$ with initial guesses for $U$ and $V$ (which can be set to zero) is conducted to determine the atomic occupation matrices required as input for the model. Subsequently, a final structural optimization using the model-predicted SC Hubbard parameters yields a self-consistent structural-electronic ground state. The computational cost of these two calculations and a model evaluation is negligible compared to the DFPT-based SC protocol they replace (see Sec. S10 in the SI). Furthermore, we have demonstrated that the full SC protocol can be reduced to just a few iterations, which prove adequate for achieving a RMSE within a few percent of the real value.
@@
Unlike other ML models designed for predicting Hubbard parameters, 27, 63, 28, 64 our model does not rely on experimental data or information from other state-of-the-art computational methods such as $GW$ and hybrid functionals, which possess inherent limitations and specific ranges of applicability. Instead, our model exclusively relies on linear-response theory through DFPT, which provides material-specific Hubbard parameters directly reflecting the local chemistry and OS of TM elements. Additionally, the current ML model can be trained on different spin configurations of the same material, with these variations reflected in the occupation matrices, though the resulting changes in Hubbard parameters are minor 83, 51. Notably, our model not only predicts on-site Hubbard $U$ parameters but also inter-site Hubbard $V$ parameters, crucial for materials characterized by significant covalent interactions, an aspect that was disregarded in previous studies. 87 Furthermore, the architecture of our ML model is highly versatile, permitting easy integration of additional inputs and outputs, facilitating exploration of diverse learning tasks beyond learning Hubbard parameters. We expect that similar hybrid ML-accelerated electronic structure methodologies, maintaining accuracy and transferability, will become prevalent, potentially yielding a comparable impact on the field as observed with ML interaction potentials.
@@
Lastly, it is essential to highlight the limitations of the trained models we have presented. These were trained on data generated using a specific computational setup (see below), which must remain exactly the same when the model is applied to other systems. In other words, the $U$ and $V$ values predicted by the model are not transferable across electronic-structure codes and even across different pseudopotentials within the same code. 88 Furthermore, the training data only encompass Fe, Mn, and Ni with Hubbard $U$ corrections, while $V$ parameters are available for pairs involving these elements and O. Consequently, these models can effectively predict Hubbard parameters for other materials with Hubbard corrections on these atoms or atom pairs, provided that the model input is generated using identical pseudopotentials, Hubbard projectors, and functional (further elaborated in the methods section). Nevertheless, with a more diverse dataset, the models could be easily extended to accommodate a broader range of compositions. Indeed, ongoing efforts aim to establish a comprehensive database of Hubbard parameters for various TM-containing materials, akin to Ref. 89, thereby significantly broadening the scope of our model to encompass numerous TM elements across diverse OS and chemical environments. While we have shown that the models can predict the results of fully self-consistent parameters, models which predict the output of single DFPT calculations may be more applicable in certain situations (see Sec. S11 in the SI). We have shown that models can be trained for this task but have not explicitly tested their performance under recursive evaluation, i.e. performing the self-consistent convergence procedure replacing DFPT completely with an ML model. The performance of the models in this use case and the potential for active learning during this process could be investigated in future work. Moreover, our model can integrate into automated AiiDA workflows, 90, 91 enabling non-experts to harness it effortlessly and access SC Hubbard parameters with minimal intervention. Consequently, we believe that our ML model represents a significant advancement in expediting materials discovery, design, and understanding based on the DFT+$U$+$V$ approach, thereby unlocking new avenues for technological progress and breakthroughs.
@@ METHODS
Ab initio calculations
All calculations are performed using the plane-wave pseudopotential method as implemented in the Quantum ESPRESSO  distribution. 92, 93, 94 We use the exchange-correlation functional constructed using $\sigma$-GGA with the PBEsol prescription, 95 and the pseudopotentials are taken from the SSSP library v1.1 (efficiency). 96, 97 For metallic ground states, we use Gaussian smearing. To construct the Hubbard projectors, we use atomic orbitals which are orthonormalized using Löwdin’s method. 98, 99 Structural optimizations are performed using DFT+$U$+$V$ 58 with the Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm 100 and convergence thresholds for the total energy of $10^{-6}$ Ry, for forces of $10^{-5}$ Ry/Bohr, and for pressure of $0.5$ Kbar. The DFPT calculations of Hubbard parameters are performed using the HP code, 101 with an accuracy of $0.01-0.1$ eV for the computed values of $U$ and $V$. The information about the kinetic-energy cutoff, $\mathbf{k}$ and $\mathbf{q}$ points sampling of the Brillouin zone for each system are detailed in Sec. S9 in the SI. More technical details can be found directly in the source files publicly available through the Materials Cloud Archive (see below).
@@
The data used to generate the results presented in this paper are accessible in the Materials Cloud Archive. 102
@@
The latest version of the code for the equivariant neural networks is available at https://github.com/camml-lab/hubbardml/. Results presented in this paper were generated using version 0.2.0 103, which is also available on the Materials Cloud Archive. 102
@@ ACKNOWLEDGEMENTS
Competing interests
The authors declare no competing interests.